OpenStudy (anonymous):

The ratio of men to women in a certain factory is 3 to 4. There are 222 men. How many workers are there? @AllTehMaffs @beccaboo333

OpenStudy (anonymous):

3:4 as a decimal is 0.75, now multiply 0.75 x 222 and you get your answer.

OpenStudy (anonymous):

incorrect \[ \frac{3}{4} = \frac{222}{x}\] x is the total number of workers

OpenStudy (anonymous):

\[x=\frac{4*222}{3}\]

OpenStudy (anonymous):

518 workers altogether

OpenStudy (anonymous):

So it's 518? ._.

OpenStudy (anonymous):

Oh never mind I see.. 4 * 222 / 3

OpenStudy (anonymous):

296?

OpenStudy (anonymous):

yup yup ^^

OpenStudy (anonymous):

Thanks again bro! :D I got more if you're able to visit the ones I tag ya in :P

OpenStudy (anonymous):

does that proportion make sense up top?

OpenStudy (anonymous):

Oh wait, I just entered the answer and it doesn't give me the correct, but it says it was wrong O_o? Thanks anyhow ;p.

OpenStudy (anonymous):

N yeah yours does, no one elses.

OpenStudy (anonymous):

It says it was wrong??

OpenStudy (anonymous):

Yeah O_o

OpenStudy (anonymous):

3/222=4/x 3x=888 x=296 women 296+222=518 total

OpenStudy (anonymous):

:/ I think @TriniEmo might have been right because I'm a dumb